Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

MGMAT test Q

Expert replies
by akshatgupta87 » Sun Apr 01, 2012 10:26 am
Q). In United States currency, a nickel is worth 5 cents, a penny is worth 1 cent, and a dime is worth 10 cents. 100 cents equals one dollar. If a hand purse contains 6 nickels, 5 pennies and 4 dimes, what is the probability of picking a coin other than a nickel twice in a row if the first coin picked is not put back?
A) 8/25
B) 12/35
C) 13/35
D) 9/25
E) 17/25

OA is B

I did the question in this way and was not able to get to the right answer:-
I found the probability of getting a nickel in the first 2 shots and subtract it from 1
i.e.
1st pick: 6/15
2nd pick: 5/14

therefore, 6*5/(15*14)
-> 1/7
therefore probability of not getting a nickel is 1- 1/7=> 6/7
But I see that I'm doing something wrong here.

So,I had to move back to the conventional method in order to get the answer
i.e.
P(D)*P(D) + P(P)*P(P) +2 P(D)P(P)

Can someone explain me, where I went wrong with my earlier approach.

~Thanks,
Akshat
Last edited by akshatgupta87 on Sun Apr 01, 2012 10:42 am, edited 1 time in total.
Join the discussion
Source: — Problem Solving |

by Bill@VeritasPrep » Sun Apr 01, 2012 10:40 am
We have 4 possible outcomes if we draw two coins:

Nickel-nickel
Nickel-something else
Something else-nickel
Something else-something else

(Since the question asks specifically about nickels, it does not matter to us if that something else is a penny or a dime. Giving us more variables than we actually need is a common probability trap; always try and simplify if possible.)

The question asks for the probability of getting something other than a nickel on both draws, so we're looking at the 4th outcome:

P(SE-SE) = 9/15 * 8/14 = 3/5 * 4/7 = 12/35
Join Veritas Prep’s 2010 Instructor of the Year, Matt Douglas for GMATT Mondays

Visit the Veritas Prep Blog

Try the FREE Veritas Prep Practice Test
Join the discussion